Find sin(2θ) Use the fact that cos(θ)=1/6 and 0°≤θ≤90°
1
Expert's answer
2011-05-30T07:43:29-0400
As sin(2θ) = 2 sinθ cosθ, and sinθ = √(1-cos2θ) =√(1 - 1/36) = √(35/36) = √35/6,

sin(2θ) = 2 * 1/6 * √35/6 = √35/18